# Canonical Autoregressive Generation

### v1

@@ Abstract
State of the art large language models are trained using large amounts of tokens derived from raw text using what is called a tokenizer. Crucially, the tokenizer determines the (token) vocabulary a model will use during inference as well as, in principle, the (token) language. This is because, while the token vocabulary may allow for different tokenizations of a string, the tokenizer always maps the string to only one of these tokenizations—the canonical tokenization. However, multiple lines of empirical evidence suggest that large language models do not always generate canonical token sequences, and this comes with several negative consequences. In this work, we first show that, to generate a canonical token sequence, a model needs to generate (partial) canonical token sequences at each step of the autoregressive generation process underpinning its functioning. Building upon this theoretical result, we introduce canonical sampling, a simple and efficient sampling method that precludes a given model from generating non-canonical token sequences. Further, we also show that, in comparison with standard sampling, the distribution of token sequences generated using canonical sampling is provably closer to the true distribution of token sequences used during training.
@@ 1 Introduction
One of the most distinctive characteristics of large language models (LLMs) is that they are trained using large amounts of text data from diverse sources, including websites, books, scientific articles and code repositories [1]. During the training process, a tokenizer breaks down the text into sequences of tokens, which are the units that make up sentences and paragraphs, and the LLM learns to predict the next token in a sequence based on the preceding tokens.
Importantly, since tokenizers typically utilize a deterministic, rule-based mapping from text to token sequences [2, 3, 4, 5], each unique string is always mapped into one tokenization referred to as the canonical tokenization [6]. Consequently, one may conclude that, at the end of the training process, LLMs will have learned to generate exclusively canonical token sequences. However, there is empirical evidence that this conclusion does not hold—LLMs do generate both canonical and non-canonical token sequences [7, 8, 6, 9, 10, 11].
In this work, we argue that the generation of non-canonical token sequences is unlikely to bring performance benefits, and it does actually have negative consequences. First, it does not increase the expressiveness of a LLM since any non-canonical token sequence can be mapped to a canonical token sequence while preserving its string-level representation. However, they can be exploited to circumvent safety guidelines because they are out-of-distribution samples [12]. Second, it can lead to a multiplicity of (plausible) tokenizations of an output string, and this creates a financial incentive for an LLM provider to strategize and misreport the (number of) tokens a model used to generate the output [11]. Third, in the context of LLM evaluation, it becomes (computationally) hard to compute the perplexity of a given string [7, 8, 6].
Motivated by the above observations, we develop a simple and efficient sampling method that precludes an LLM from generating non-canonical token sequences. Our method is based on the following theoretical result, which may be of independent interest: to generate a canonical token sequence, a model needs to generate (partial) canonical token sequences at each step of the autoregressive generation process underpinning its functioning. More specifically, building upon this result, our method utilizes the Gumbel-Max trick to efficiently sample only from the subset of tokens that can lead to (partial) canonical token sequences at each step of the generation. Importantly, we are able to show that, in comparison with standard sampling, the distribution of token sequences generated by canonical sampling is provably closer to the true distribution of token sequences used during training.
Related work. The study of tokenization has a rich history in natural language processing [13, 14]. More recently, in the context of LLMs, there has been a renewed interest in formalizing tokenization and analyzing its properties [15, 16, 17], with the Byte-Pair Encoding (BPE) tokenization algorithm in particular receiving increased attention [18, 19, 20]. The impact of tokenization on LLMs has also been studied empirically [21, 22] in generation tasks involving foreign languages [23], translation [24], arithmetic [25], mental health [26], and privacy [27, 28], among others. Moreover, there have also been efforts to design token-free LLMs that operate at the character or byte level [29, 30, 31, 32, 33], as well as stochastic tokenizers [4, 34]. However, none of the above works has paid attention to the existence and significance of non-canonical tokenizations.
Only very recently, there has been a paucity of work studying the impact of non-canonical tokenizations on text perplexity calculations [7, 8, 6, 9, 10], safety guidelines [12], and financial incentives of LLM providers [11]. However, none of the above works has studied how to prevent the generation of non-canonical token sequences, which is the focus of our work.
@@ 2 Preliminaries
In this section, we first define and formally characterize (deterministic) tokenizers and canonical tokenizations. Then, we briefly review the popular Byte-Pair Encoding (BPE) tokenization algorithm [2, 3]. Finally, we conclude with a description of the aspects of LLM training and generation that are relevant for our work.
@@
Tokenizers and canonical tokenizations. Tokenizers are tools that operate on sequences of characters (i.e., strings) and sequences of tokens, and can transform one type into the other. Formally, let $\Sigma$ be a finite set of characters and $\Sigma^{*}$ be the set of all finite strings that can be created using the characters in $\Sigma$. Similarly, let $V$ be a finite set of tokens, which we will refer to as the vocabulary, and $V^{*}$ be the set of all finite token sequences that can be created using the tokens in $V$. Then, a tokenizer ${\mathcal{T}}$ is characterized by a tuple ${\mathcal{T}}\coloneqq(\Sigma,V,\texttt{enc},\texttt{dec})$, where $\texttt{enc}:\Sigma^{*}\rightarrow V^{*}$ is an encoder, which transforms strings to token sequences, and $\texttt{dec}:V^{*}\rightarrow\Sigma^{*}$ is a decoder, which transforms token sequences to strings.
@@
Let $\bm{\sigma}$ be a string and $\mathbf{s}\in V^{*}$ be a token sequence such that $\texttt{dec}(\mathbf{s})=\bm{\sigma}$, then, we will say that $\mathbf{s}$ is a (valid) tokenization of the string $\bm{\sigma}$. Here, note that there may be multiple tokenizations of a string $\bm{\sigma}$, that is, there may exist $\mathbf{s},\mathbf{s}^{\prime}\in V^{*}$ such that $\mathbf{s}\neq\mathbf{s}^{\prime}$ and $\texttt{dec}(\mathbf{s})=\texttt{dec}(\mathbf{s}^{\prime})=\bm{\sigma}$. However, given a string $\bm{\sigma}$, the encoder deterministically picks a single tokenization $\texttt{enc}(\bm{\sigma})$ among all tokenizations of $\bm{\sigma}$, which is often called the canonical tokenization [6].
The BPE tokenization algorithm. There exist many tokenization algorithms to construct the set of characters $\Sigma$, the set of tokens $V$, the encoder enc, and the decoder dec characterizing a tokenizer ${\mathcal{T}}$. Here, we focus on the BPE tokenization algorithm [2, 3], since it is the algorithm used by most, if not all, state-of-the-art LLMs. [^1] In Appendix B, we review Unigram [4] and Wordpiece [5], which are also well-known tokenization algorithms. In a nutshell, the BPE algorithm aims to create a tokenizer ${\mathcal{T}}$ with a set of tokens $V$ corresponding to character sequences that appear frequently in a training set of strings $\mathcal{C}$. To this end, it proceeds as follows.
In an initialization phase, the algorithm sets $\Sigma$ to be the set of all characters that appear at least once in $\mathcal{C}$, $V$ to be the set of single-character tokens, that is, for each $c\in\Sigma$, there exists one and only one $t\in V$ such that $\texttt{dec}(t)=c$, and ${\mathcal{S}}$ to be the set of single-character token sequences $\mathbf{s}\in V^{*}$ representing all strings in $\mathcal{C}$. After the initialization phase, the algorithm proceeds iteratively for a predetermined number of iterations. At each iteration, it looks for the pair of tokens $t,t^{\prime}\in V$ whose concatenation $t\shortmid t^{\prime}$ appears most frequently in the set of token sequences ${\mathcal{S}}$, it creates a new token $t\circ t^{\prime}$, where the symbol $\circ$ denotes the merge operation and $\texttt{dec}(t\circ t^{\prime})=\texttt{dec}(t)\shortmid\texttt{dec}(t^{\prime})$, and it adds the newly created token to $V$. Then, for each token sequence $\mathbf{s}\in{\mathcal{S}}$, it replaces all occurrences of $t\shortmid t^{\prime}$ by $t\circ t^{\prime}$ one by one. Lastly, it creates a merge rule $r_{t,t^{\prime}}$, which specifies the replacement of $t\shortmid t^{\prime}$ with $t\circ t^{\prime}$, and adds it to an ordered sequence of merge rules $\mathcal{R}$.
After termination, the algorithm defines the encoder enc and decoder dec as follows. For any given token sequence $\mathbf{s}\in V^{*}$, $\texttt{dec}(\mathbf{s})$ transforms the sequence to a string one token at a time, in order, using the token definitions. For any given string $\bm{\sigma}\in\Sigma^{*}$, $\texttt{enc}(\bm{\sigma})$ first transforms the string to a sequence of single-character tokens. Then, it merges consecutive tokens from this sequence following the merge rules from $\mathcal{R}$, in order, until no merge rule is applicable, and it returns the resulting sequence—the canonical sequence. [^2] If $t\shortmid t^{\prime}$ appears multiple times in a token sequence, the merge rule $r_{t,t^{\prime}}$ is applied in order of appearance in the sequence.
@@
LLM training and generation. During training, an LLM learns to predict the next-token in canonical sequences of tokens derived from raw text using a tokenizer. More formally, given a (partial) token sequence $\mathbf{s}\in V^{*}$, the goal is (typically) to minimize the (cross-entropy) loss between the model’s predicted distribution $d_{\mathbf{s}}\in\Delta(V)$ and the true next-token distribution $p_{\mathbf{s}}=P[T\mid\bm{S}=\mathbf{s}]$.
@@
During generation, an LLM takes as input a prompt sequence $s_{q}\in V^{*}$ and responds with an output sequence $s\in V^{*}$, generated using an autoregressive process. At each time step of the process, the LLM first takes as input the concatenation of the prompt sequence $s_{q}$ and the (partial) output sequence $s$, and generates a distribution over tokens $d_{s_{q}\shortmid s}\in\Delta(V)$. Then, it samples the next token $t$ from the distribution $d_{s_{q}\shortmid s}$ and appends the token $t$ to the output sequence $s$. The process continues until a special end-of-sequence token is sampled.
@@
Importantly, since LLMs are trained on finite data, the support of the distribution $d_{\mathbf{s}_{q}\shortmid\mathbf{s}}$ may differ from the respective true distribution $p_{\mathbf{s}_{q}\shortmid\mathbf{s}}$. As a result, it is possible for an LLM to generate a non-canonical token sequence, even if it has encountered no such sequences during training [7, 8, 6, 9, 10, 12, 11].
3 An Efficient Approach to Canonical Autoregressive Generation

### v2

# Tokenization Multiplicity Leads to Arbitrary Price Variation in LLM-as-a-service

@@ Abstract
Providers of LLM-as-a-service have predominantly adopted a simple pricing model: users pay a fixed price per token. Consequently, one may think that the price two different users would pay for the same output string under the same input prompt is the same. In our work, we show that, surprisingly, this is not (always) true. We find empirical evidence that, particularly for non-english outputs, both proprietary and open-weights LLMs often generate the same (output) string with multiple different tokenizations, even under the same input prompt, and this in turn leads to arbitrary price variation. To address the problem of tokenization multiplicity, we introduce canonical generation, a type of constrained generation that restricts LLMs to only generate canonical tokenizations—the unique tokenization in which each string is tokenized during the training process of an LLM. Further, we introduce an efficient sampling algorithm for canonical generation based on the Gumbel-Max trick. Experiments on a variety of natural language tasks demonstrate that canonical generation is comparable to standard generation in terms of performance and runtime, and it solves the problem of tokenization multiplicity.
@@ 1 Introduction
Imagine you run an online service that offers AI-powered translation to help tourists seamlessly navigate websites in languages other than their mother tongue. Behind the scenes, for each website the users visit, your service sends the website’s text to a provider of an LLM and asks their model to generate the translation in the mother tongue of the user. One day, reviewing your costs, you notice something strange: Hans and Emma, two users of your service, visited the same page on a website with geographical fun facts, requesting a German translation of the sentence
“The Acari River is a river of Minas Gerais state in southeastern Brazil”.
Both of them received the same translated text down to the last character, yet, you were charged different amounts for each by the LLM provider. How can identical results have different costs?
The computational resources required to operate state-of-the-art large language models are too large for most (enterprise) users to run them locally (Narayanan et al., 2021; Jiang et al., 2024). As a consequence, users are relying on a growing market of cloud-based providers who offer access to such models via an application programming interface (API) (Sun et al., 2022; La Malfa et al., 2024; Pais et al., 2022; Liagkou et al., 2024). The way that providers of LLM-as-a-service have come to charge the use of their models is heavily dependent on one of the most distinctive technical characteristics shared by most, if not all, state-of-the-art models—they process and generate information in discrete units called tokens. [^1] Tokens are (sub-)words, symbols and numbers that make up sentences and paragraphs. In particular, since the computational resources required to process an input prompt and generate a response are directly proportional to the number of tokens involved (Samsi et al., 2023; Fernandez et al., 2025), providers have predominantly adopted a straightforward pay-per-token pricing model—users pay a fixed price per token.
Under the pay-per-token pricing model, one may naturally assume that, if i) two users submit the same prompt to a provider, ii) the provider feeds the prompt into the same model, and iii) both users receive exactly the same output string, then, they will pay the same price. In our work, we show that, surprisingly, this assumption does not hold. We find empirical evidence that, particularly for non-english outputs, both proprietary and open-weights LLMs often generate the same (output) string with different tokenizations, even under the same input prompt, and this multiplicity of tokenizations in turn leads to price variation. Moreover, since users derive value from the text that an output token sequence represents, rather than the token sequence itself, we argue that this price variation is arbitrary and undesirable.
To address the problem of tokenization multiplicity, we introduce canonical generation, a type of constrained generation that restricts LLMs to only generate canonical tokenizations—the unique tokenization in which each string is tokenized during the training process of an LLM (Geh et al., 2024). In doing so, we also introduce an efficient sampling algorithm for canonical generation based on the Gumbel-Max trick that leverages the following theoretical result, which may be of independent interest: to generate a canonical tokenization, a model needs to generate (partial) canonical tokenizations at each step of the generation process underpinning its functioning.
In addition to solving the problem of tokenization multiplicity, we show that, in comparison with standard generation, the distribution of token sequences generated by canonical generation is provably closer to the true distribution of token sequences used during training and, in practice, the performance and runtime of canonical generation are comparable to standard generation. We have released all code and data used in our experiments at: https://github.com/Networks-Learning/Tokenization-Multiplicity.
Further related work. Our work builds upon further related work on tokenization in LLMs, multilingual LLMs, and the economics of LLM-as-a-service.
The study of tokenization has a rich history in natural language processing (Palmer, 2000; Jurafsky and Martin, 2025). More recently, in the context of LLMs, there has been a renewed interest in formalizing tokenization and analyzing its properties (Gastaldi et al., 2025; Phan et al., 2024; Rajaraman et al., 2025), with the Byte-Pair Encoding (BPE) tokenization algorithm in particular receiving increased attention (Berglund and van der Merwe, 2023; Zouhar et al., 2023; Kozma and Voderholzer, 2024). The impact of tokenization on LLMs has also been studied empirically (Hou et al., 2023; Athiwaratkun et al., 2024) in generation tasks involving foreign languages (Fujii et al., 2023), translation (Domingo et al., 2019), arithmetic (Singh and Strouse, 2024), mental health (Liu et al., 2023), and privacy (Kharitonov et al., 2021; Petrov et al., 2023), among others. Moreover, there is a recent line of work studying the impact of non-canonical tokenizations on text perplexity calculations (Cao and Rimell, 2021; Chirkova et al., 2023; Geh et al., 2024; Vieira et al., 2025a; Giulianelli et al., 2024), safety guidelines (Geh et al., 2025), downstream tasks (Zheng et al., 2025), and image watermarking (Jovanović et al., 2025), as well as an effort to circumvent non-canonical tokenizations occuring as a consequence of unusual input prompt endings (Guidance, 2023). Within this line of work, the work most closely related to ours is by Vieira et al. (2025b), who have also independently proposed canonical generation [^2] A preliminary version of our work and their work were posted on arXiv within three days of each other.. However, their work does not provide empirical evidence of tokenization multiplicity and its influence on the pay-per-token pricing model as we do, their sampling algorithms for canonical generation are computationally less efficient compared to ours, and their experimental evaluation does not demonstrate that canonical generation is comparable to standard generation in terms of performance across natural language tasks.
There has been extensive research on the capabilities of multilingual LLMs, including their performance on downstream tasks (Fujii et al., 2023; Zhang et al., 2022; Rust et al., 2021) as well as their safety vulnerabilities (Shen et al., 2024; Wang et al., 2024; Deng et al., 2024; Dong et al., 2025) across different languages. Within this research, the work most related to ours is by Ahia et al. (2023), who have shown that multilingual LLMs require more tokens to generate text of similar meaning in minority languages than in english. Our work complements their work by providing empirical evidence that multilingual LLMs suffer from tokenization multiplicity in minority languages.
The rapidly growing literature on the economics of LLM-as-a-service (Mahmood, 2024; Laufer et al., 2024; Cai et al., 2025; Saig et al., 2024; Bergemann et al., 2025; Sun et al., 2025a, b; Artola Velasco et al., 2025; Velasco et al., 2026) has predominantly focused on understanding the incentives providers may have to act strategically at the expense of users. Within this literature, the work most closely related to ours is by Artola Velasco et al. (2025), who show that token multiplicity enables an unfaithful provider to strategize and misreport the tokenization of an output generated by a model they serve without raising suspicion. Moreover, to eliminate the incentive to strategize, they propose an incentive-compatible pay-per-character pricing model, which in turn also eliminates the price variation due to tokenization multiplicity. However, in their work, they do not empirically demonstrate that token multiplicity can occur in practice, and it can lead to arbitrary price variation even under faithful providers, as we do in our work. Further, it is worth noting that the pay-per-character pricing model makes a provider’s profit margin vary across tokens, and this may discourage its practical adoption.
@@ 2 Preliminaries
In this section, we first define and formally characterize (deterministic) tokenizers and canonical tokenizations. Then, we briefly review the aspects of LLM training and generation that are relevant for our work.
@@
Tokenizers and canonical tokenizations. Tokenizers are tools that operate on sequences of characters (i.e., strings) and sequences of tokens, and can transform one type into the other. Formally, let $\Sigma$ be a finite set of characters and $\Sigma^{+}$ be the set of all finite strings that can be created using the characters in $\Sigma$. Similarly, let $V$ be a finite set of tokens, which we will refer to as the vocabulary, and $V^{+}$ be the set of all finite token sequences that can be created using the tokens in $V$. Then, a tokenizer $\mathcal{T}$} is characterized by a tuple ${\mathcal{T}}\coloneq(\Sigma,V,\texttt{enc},\texttt{dec})$, where $\texttt{enc}:\Sigma^{+}\rightarrow V^{+}$ is an encoder, which transforms strings to token sequences, and $\texttt{dec}:V^{+}\rightarrow\Sigma^{+}$ is a decoder, which transforms token sequences to strings.
@@
Let $\bm{\sigma}$ be a string and $\mathbf{s}\in V^{+}$ be a token sequence such that $\texttt{dec}(\mathbf{s})=\bm{\sigma}$, then, we will say that $\mathbf{s}$ is a (valid) tokenization of the string $\bm{\sigma}$. Here, note that there may be multiple tokenizations of a string $\bm{\sigma}$, that is, there may exist $\mathbf{s},\mathbf{s}^{\prime}\in V^{+}$ such that $\mathbf{s}\neq\mathbf{s}^{\prime}$ and $\texttt{dec}(\mathbf{s})=\texttt{dec}(\mathbf{s}^{\prime})=\bm{\sigma}$. However, given a string $\bm{\sigma}$, the encoder deterministically picks a single tokenization $\texttt{enc}(\bm{\sigma})$ among all tokenizations of $\bm{\sigma}$, which is often called the canonical tokenization (Geh et al., 2024). For details on the most commonly used tokenization algorithms—BPE (Gage, 1994; Sennrich et al., 2016), Unigram (Kudo, 2018) and Wordpiece (Song et al., 2021)—refer to Appendix A.
@@
LLM training and generation. During training, an LLM learns to predict the next token in canonical sequences of tokens derived from raw text using a tokenizer. More formally, let $p_{\mathbf{s}}=P[T\mid\bm{S}=\mathbf{s}]$ denote the true distribution of the random variable $T\in V$, representing the next token given a (partial) token sequence $\mathbf{s}\in V^{+}$. Then, the goal of LLM training is (typically) to minimize the (cross-entropy) loss between the model’s predicted distribution $d_{\mathbf{s}}\in\Delta(V)$ and the true next-token distribution $p_{\mathbf{s}}$.
@@
During generation, an LLM takes as input a prompt sequence $\mathbf{s}_{q}\in V^{+}$ and responds with an output sequence $\mathbf{s}\in V^{+}$, generated using an autoregressive process. At each time step of the process, the LLM first takes as input the concatenation of the prompt sequence $\mathbf{s}_{q}$ and the (partial) output sequence $\mathbf{s}$, and generates a distribution over tokens $d_{\mathbf{s}_{q}\shortmid\mathbf{s}}\in\Delta(V)$. Then, it samples the next token $t$ from the distribution $d_{\mathbf{s}_{q}\shortmid\mathbf{s}}$ and appends the token $t$ to the output sequence $\mathbf{s}$. The process continues until a special end-of-sequence token is sampled. For the remainder of this paper, we will omit the prompt sequence $\mathbf{s}_{q}$ from the notation and write $d_{\mathbf{s}}$ and $p_{\mathbf{s}}$ for brevity.
@@
Importantly, since LLMs are trained on finite data, the support of the distribution $d_{\mathbf{s}}$ may differ from the respective true distribution $p_{\mathbf{s}}$. As a result, it is possible for an LLM to generate a non-canonical token sequence, even if it has encountered no such sequences during training (Cao and Rimell, 2021; Chirkova et al., 2023; Vieira et al., 2025a; Giulianelli et al., 2024; Geh et al., 2025).
3 Can Hans and Emma Receive Different Tokenizations for the Same Output String?
We answer this question affirmatively. Using three simple natural language tasks, we demonstrate that tokenization multiplicity can occur in both proprietary and open-weights LLMs. Specifically, we focus on translation (as in the example of Hans and Emma in Section 1), spell checking, and rephrasing. The motivation for these choices is that, in all these tasks, an LLM has to process and rewrite a given input text and is therefore more likely to generate the same output string despite the randomness in its generation.
Figure 1: Translation task example. The top box shows the input prompt, which consists of a translation instruction and the accompanying Wikipedia text to be translated. The latter two boxes show two outputs generated by Qwen2.5-7B-Instruct as response to the input prompt, corresponding to the same string but with two different tokenizations.
For each of the three tasks, we construct $100$ input prompts using short texts from Wikipedia. [^3] https://dumps.wikimedia.org/ For the translation task, each prompt consists of a (system) instruction to translate a short text written in English to one of $5$ target languages, followed by the respective Wikipedia text in the English language. For the other two tasks, each prompt consists of a (system) instruction to spell check or rephrase short texts written in $6$ different languages (including English) and, for the spell checking task, we first introduce a small number of typos in each Wikipedia text by replacing a few latin characters with other latin characters selected at random. To simulate scenarios where different users ask an LLM to perform the same task on the same input text, we feed each input prompt to the LLM $100$ times, keeping all parameters identical except for the random seed used during generation. We then focus on (the existence of) pairs of outputs whose strings are identical but their tokenization lengths differ.
As a starting point, in Figure 1, we illustrate the translation task that we consider in our experiments using as an example a pair of outputs generated by Qwen2.5-7B-Instruct (Yang et al., 2024) to the input text introduced at the beginning of Section 1. We observe that, while both output pairs correspond to the same string, Hans’s output consists of $26$ tokens while Emma’s output consists of $28$ tokens due to a difference in the generated tokenization of the words “südöstlichen” and “Brasilien”. Importantly, these tokenization differences cause the latter output to be $7.7\%$ more expensive than the former under pay-per-token pricing. For similar examples illustrating the spell checking and rephrasing tasks, refer to Appendix E.
Further, we proceed to quantify the frequency and magnitude of such inconsistencies in pricing across different tasks, LLMs, and languages. In our experiments, we consider both proprietary LLMs from the gpt, gemini, and claude families and open-weights LLMs from the Llama and Qwen families. For brevity, in the remainder of the section, we use shortened names to refer to each model, and we provide the full names of all models in Table 3 in Appendix F, along with additional details about the hardware, data, and APIs we use. [^4] For (proprietary) gpt, gemini, and claude models, we used the official LLM-as-a-service APIs from OpenAI, Google, and Anthropic, respectively. For open-weights models, we ran all experiments locally, however, note that a user without access to specialized hardware would normally also access them via LLM-as-a-service APIs from third-party providers.
We first look into the (conditional) probability that, once two users receive the same output string from the model, the lengths of the two tokenizations differ. Formally, let $\mathbf{S}$ and $\mathbf{S}^{\prime}$ denote the random variables corresponding to tokenizations received by two different users as a response after providing the same prompt, i.e., $Q=Q^{\prime}$. Our goal is to estimate the quantity $P(\texttt{len}(\mathbf{S})\neq\texttt{len}(\mathbf{S}^{\prime})\mid\texttt{dec}(\mathbf{S})=\texttt{dec}(\mathbf{S}^{\prime}),Q=Q^{\prime})$. To this end, for each prompt, we count the number of output pairs whose strings match but tokenization lengths differ as a fraction of all output pairs whose strings match, and we take the average across prompts. Note that, to perform fair comparisons across models, we intentionally focus on cases where there are differences in the tokenization length rather than the tokenization itself, since the generated tokenization is not observable under all APIs. [^5] The API services for gpt5m, gemini and claude return only the output string without disclosing the tokenization that was generated. In those cases, it is still possible to identify differences in tokenization length, which is always disclosed since pricing directly depends on it. Figure 2 summarizes the results for German language, which show that tokenization multiplicity can indeed occur both when using open-weights and proprietary models. In particular, for all models except gemini, we find cases of tokenization multiplicity in at least one task and, for open-weights models (i.e., Llama8b and Qwen7b), we find that tokenization multiplicity occurs regularly across all three tasks. Refer to Appendix G.1 for qualitatively similar results in other languages. In this context, note that the gemini model did generate a few (<<11%) non-canonical outputs, which suggests that this model may also present tokenization multiplicity, but we did not observe it due to the finite sample size of our experiments. [^6] Refer to Appendix G.1 for additional results showing the percentage of non-canonical outputs across all models and tasks.
Figure 2: Probability of tokenization multiplicity. The plots show the empirical probability that the length of two output tokenizations to the same input prompt differ, conditioned on the output strings being the same. Each panel corresponds to one of the three tasks we consider in our experiments involving outputs in the German language. Across all panels, error bars represent $95\%$ confidence intervals resulting from $100$ input prompts.
Next, we focus on outputs where tokenization multiplicity does occur and look into the magnitude of the variation in tokenization lengths, which directly determines the variation in prices charged to users under pay-per-token pricing. Specifically, for each output string $\bm{\sigma}$ where there are at least two tokenizations with different length, we measure the relative difference in length (and hence price) between the shortest tokenization $\mathbf{s}_{min}$ and the longest tokenization $\mathbf{s}_{max}$, i.e., $\left(\texttt{len}\left(\mathbf{s}_{max}\right)-\texttt{len}\left(\mathbf{s}_{min}\right)\right)/\texttt{len}\left(\mathbf{s}_{min}\right)$. Figure 3 summarizes the results for German language, which show that, whenever tokenization multiplicity occurs, a user may pay up to $15\%$ higher price in comparison with another user for the same output string, with similar levels of variation across both open-weights and proprietary models. Refer to Appendix G.1 for qualitatively similar results in other languages.
Further, we analyze how the prevalence of tokenization multiplicity changes depending on the language of the output. To this end, out of the $100$ input prompts we construct per language and task, we measure the number of prompts for which a model (here, gpt5m) generates at least two outputs with the same string but different tokenization lengths. Figure 4 summarizes the results, which show that, across all three tasks, tokenization multiplicity is more prevalent in minority languages. For example, even though ChatGPT officially supports languages such as Turkish or Swahili, [^7] https://help.openai.com/en/articles/8357869 we observed that at least $7\%$ of the input prompts for these two languages led gpt5m to generate identical outputs with different prices. For qualitatively similar results with other models, refer to Appendix G.1.
Figure 3: Magnitude of price variation. The plots show the empirical distribution of the relative difference in length between the longest and shortest tokenization of each output string, across all outputs where tokenization multiplicity occurs. Each panel corresponds to one of the three tasks we consider in our experiments involving outputs in the German language. Across all panels, box plots show the quartiles of the respective distributions with black horizontal lines representing median values.
To study the prevalence and magnitude of tokenization multiplicity, we have focused so far on pairs of short outputs corresponding to exactly the same string, since this allows us to attribute any price variation exclusively to token multiplicity. In this context, one may think that, on pairs of longer outputs with multiple partial string matches, tokenization multiplicity may not lead to significant price variation if the difference in tokenization length across partial matches cancels out. However, in what follows, we provide empirical evidence that the difference in tokenization length is not independent across partial matches and, as a consequence, the price variation due to token multiplicity may be more pronounced in longer outputs. In particular, in outputs generated by gpt4om to the translation task from English to German, but using longer Wikipedia texts (refer to Appendix F for details regarding the experimental setup), we find that, if a word is generated with a non-canonical tokenization, then, subsequent occurrences of the word are typically generated using that same tokenization. More specifically, we observe that, in $88\%$ of those cases, all subsequent occurrences (up to $10$) are generated using the same non-canonical tokenization. In contrast, in $10\%$ of the remaining cases, all subsequent occurrences of the word are generated using the canonical tokenization and, in the last $2\%$ of cases, subsequent occurrences use a mix of both canonical and non-canonical tokenizations. As an immediate consequence, it is easy to find examples of long, similar outputs containing multiple repetitions of the same word with different but consistent tokenizations, as shown in Figure 6 in Appendix E.
4 Avoiding Tokenization Multiplicity through Canonical Generation
To solve the problem of tokenization multiplicity, we introduce canonical generation, a type of constrained generation that restricts LLMs to generate each output string with only its canonical tokenization. In this context, we argue that, if one were to pick a single tokenization for each output string, the canonical tokenization presents itself as the most natural choice. This is because, during the training of an LLM, all strings in the training data are first encoded canonically using a tokenizer, and the LLM is then trained to complete these canonical token sequences.